find the equation of the line containing the given pair (-5,-3) (-2,-4).
I came up with:
y=1/3x-4/3 but the systems says that it is not correct.
Please help! Thank you

Answer by scott8148(6628)

the slope (m) is -1/3 (not 1/3)

substituting to find the y-intercept (b) __ -3=(-1/3)(-5)+b __ -3=(5/3)+b __ -14/3=b

y=(-1/3)x-(14/3)